OpenStudy (anonymous):

How do I find if its even or odd with a cubed root like: f(x)=cubed root of 6x^2

OpenStudy (zarkon):

this \[\sqrt[3]{6x^2}\]?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

even

OpenStudy (anonymous):

how do I find that out though

OpenStudy (zarkon):

\[f(-x)=\sqrt[3]{6(-x)^2}=\sqrt[3]{6x^2}=f(x)\]